# PointAD+: Learning Hierarchical Representations for ZS 3D Anomaly Detection

### v4

@@ 3.6.1 Implicit point representation learning on the rendering layer (PointAD)
$\hskip-2.84544pt\textstyle L_{I(3d)}^{global}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}).$ (6)
@@ 3.6.3 Cross-hierarchy contrastive alignment
$-\log\frac{\exp(\cos(g^{g}_{n},g^{r}_{n}))}{\exp(\cos(g^{g}_{n},g^{r}_{n}))+\exp(\cos(g^{g}_{n},g^{r}_{a}))}$ (18)
$-\log\frac{\exp(\cos(g^{r}_{a},g^{g}_{a}))}{\exp(\cos(g^{r}_{a},g^{g}_{a}))+\exp(\cos(g^{r}_{a},g^{g}_{n}))}$
$-\log\frac{\exp(\cos(g^{r}_{n},g^{g}_{n}))}{\exp(\cos(g^{r}_{n},g^{g}_{n}))+\exp(\cos(g^{r}_{n},g^{g}_{a}))}$
@@ 4.1 Experiment Setup
Dataset. We evaluate the performance of ZS 3D anomaly detection on three publicly available 3D anomaly detection datasets, MVTec3D-AD, Eyecandies, and Real3D-AD. MVTec3D-AD comprises 4147 point clouds across 10 categories. These objects exhibit diverse object semantics, including bagel, cable gland, carrot, cookie, dowel, foam, peach, potato, rope, and tire. The training dataset comprises 2656 normal point clouds, and the test dataset includes 948 normal and 249 anomaly point clouds, covering several anomaly types. Point-wise annotations are available for the point clouds. MVTec3D-AD also provides corresponding 2D-RGB image counterparts for the point clouds. We remove the background plane of point clouds in the whole dataset like [19]. Eyecandies also has 10 different classes and provides the corresponding 2D RGB information. The more recent Real3D-AD is a challenging dataset containing stereo point clouds with 12 categories, without the corresponding RGB information.
@@
Implementation details. Both point clouds and 2D renderings are resized to 336 $\times$ 336. We use Open3D library to generate 9 views by rotating point clouds along the X-axis at angles of $\{-\frac{4}{5}\pi,-\frac{3}{5}\pi,-\frac{2}{5}\pi,-\frac{1}{5}\pi,0,\frac{1}{5}\pi,\frac{2}{5}\pi,\frac{3}{5}\pi,\frac{4}{5}\pi\}$ for most categories. Some objects lose their surface when rendered at certain angles, we manually adjust them to an appropriate angle. The backbone of PointAD+ and PointAD is the pre-trained CLIP model (VIT-L/14@336px in open_clip). Following [16], we improve the local visual semantics of the vision encoder of CLIP without modifying its parameters. During training, we keep all parameters of CLIP frozen and set the learnable word embeddings in object-agnostic text templates to $12$. The hyperparameters, including the number of neighbors $k$, the weighting scalar $\alpha$, and the scale parameter $\sigma$ of the Gaussian kernel, are set to 10, 0.5, and 1, respectively. We use the Adam optimizer with a learning rate of 0.001 to optimize the learnable parameters. The experiment runs for 15 epochs with a batch size of 4. All experiments were conducted on a single NVIDIA RTX 3090 24GB GPU using PyTorch-2.0.0. PointAD+ and PointAD by default use object-agnostic text prompts, whereas PointAD+-CoOp and PointAD-CoOp employ object-aware prompts.
Baselines. As ZS 3D and multimodal 3D anomaly detection remain underexplored areas in the field, we make a great effort to provide these comparisons. We apply the original CLIP to our framework for 3D detection, called CLIP + Rendering. Also, we reproduce SOTA 3D recognition works including PointCLIP V2 [54] and Cheraghian [55] and adapt them for ZS 3D anomaly detection. We compare the SOTA 2D anomaly detection approach AnomalyCLIP [16] by fine-tuning it on depth maps. All hyperparameters in these baselines are kept the same. We will present the detailed reproduction:
  • CLIP + Rendering is a method, where we apply the original CLIP into our framework for ZS 3D anomaly detection. It uses the same rendering procedure as PointAD+ and PointAD. Following [11, 16], we integrate anomaly semantics into CLIP by two class text prompt templates: A photo of a normal [cls] and A photo of an anomalous [cls], where cls denotes the target class name.
@@
  • Cheraghian (IJCV 2022) is an approach for ZS 3D classification without foundation models. They directly extract the point presentations by PointNet and use word2vector [56] to generate the textual embedding of [c]. To incorporate the anomaly semantics into Cheraghian, we average the textual embeddings of [c] and damaged. We replace the global representation with dense representations to provide the segmentation.
@@ 4.2 Main Results
ZS 3D anomaly detection. Table I presents the comparison of ZS 3D performance. Compared to the point-based method Cheraghian and the projection-based method PointCLIP V2, PointAD+ and PointAD achieve superior performance on ZS 3D anomaly detection over all three datasets across 32 object semantics. Especially, PointAD outperforms CLIP + Rendering from 61.2% to 82.0% I-AUROC and from 85.8% to 94.2% AP on MVTec3D-AD. PointAD+ further obtains the performance gain over PointAD. Besides global anomaly detection, PointAD+ and PointAD also attain superior segmentation performance. PointAD improves MVTec3D-AD by a large margin compared to Cheraghian from 88.2% to 95.5% P-AUROC and from 57.0% to 84.4% AUPRO. This improvement in overall performance is attributed to PointAD adapting CLIP’s strong generalization to glocal anomaly semantics through hybrid representation learning. On this basis, PointAD+ further improves the performance from 95.5% to 96.8% P-AUROC and from 84.4% to 88.4% AUPRO. It is attributed to the hierarchy representation learning, incorporating explicit point relations to better reveal the abnormality within point clouds. As shown in Figure 5, PointAD+ and PointAD could accurately segment the unseen anomalies across diverse objects. In addition, PointAD+ and PointAD advance PointAD+-CoOp and PointAD-CoOp across all datasets by blocking the class semantics in text prompts [16].
ZS multimodal 3D anomaly detection. PointAD+ and PointAD allow the direct incorporation of RGB counterpart when it is available during the test. Table II indicates that PointAD+ and PointAD leverage additional RGB information in a plug-and-plug manner to boost their performance. PointAD gets the performance gain from 82.0% to 86.9% AUROC and from 94.2% to 96.1% AP for global semantics on MVTec3D-AD. Additionally, as for local semantics, the performance improves from 95.5% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. A large performance gain is also obtained on Eyecandies and Real3D-AD. PointAD+ can also benefit from the RGB information and break the performance achieved by PointAD, i.e., 98.2% vs. 97.2% P-AUROC and 93.1% vs. 90.2% AUPRO on MVTec3D-AD, and 79.2% vs. 77.7% I-AUROC and 81.7% vs. 80.4% AP on Eyecandies. As other methods can improve their performance in some metrics, they still suffer from performance degradation in other metrics due to inefficient integration of the two modalities. Instead, PointAD+ achieves overall improvement by incorporating explicit joint constraints on both point and pixel information.
@@
4.3 Result Analysis
PointAD+ characterizes itself by capturing abnormality from two aspects: implicit 3D abnormality where point anomalies are projected into the pixel space; explicit 3D abnormality, which reveals the inherent point relations. In this section, we provide a more in-depth analysis of how they contribute to the success of PointAD+.
4.3.1 Implicit point learning on rendering layer
We first investigate how rendering anomalies facilitate 3D anomaly detection. We visualize the anomaly score maps of the 3D and corresponding 2D counterparts of PointAD+ on MVTec3D-AD. And, we also visualize the anomaly score map of 2D counterparts according to Eq. 9, projected by 3D point anomalies. It can be observed that PointAD+ has a strong detection ability on such 2D anomalies. The strong representative pixel representations from multiple views facilitate more precise 3D anomaly detection. We conduct a quantitative analysis of the view number in the ablation study. The strong 3D and 2D detection capabilities of PointAD+ are from hybrid representation learning, which not only enables PointAD+ to capture the 3D anomalies but also explicitly constrains 2D representations.
4.3.2 Explicit point learning on geometric layer
We further examine what PointAD+ learns from the geometry layer according to Eq. 15. As shown in Figure LABEL:fig:_ei1, explicit point learning effectively captures point-level abnormalities in previously unseen objects. By incorporating geometric information, it demonstrates a stronger ability to detect abnormal point relations, making it more accurate than implicit point learning in identifying such anomalies. However, in Figure LABEL:fig:_ei2, when point relations change rapidly, explicit point representations may produce false anomaly alarms.
@@ 4.3.2 Explicit point learning on geometric layer
TABLE V: Analysis on aggregation strategy $\ell$.
$\ell$ Point detection Multimodal detection
Local Global Local Global
Mean (96.9, 88.7) (82.5, 94.8) (98.2, 93.1) (87.7, 96.6)
Max (97.0, 88.3) (82.5, 94.4) (97.9, 93.5) (87.8, 96.4)
TABLE VI: Comparison of computation overhead with SOTA approaches on MVTec3D-AD.
@@
PointCLIP V2 (ZS) 4660.49 0.66 9747MB (87.4, 52.3) (78.3, 49.4)
CLIP + Rendering (ZS) 1057.36 3.61 3685MB (61.2, 85.8) (-, 54.4)
Cheraghian (ZS) 31.49 2.86 4847MB (88.2, 57.0) (53.6, 81.7)
AnomalyCLIP (ZS) 192.41 5.26 3348MB (88.9, 60.9) (56.4, 83.5)
PointAD (ZS) 1656.34 2.52 4275MB (95.5, 84.4) (82.0, 94.2)
PointAD+ (ZS) 1658.34 2.18 5811MB (96.9, 88.7) (82.5, 94.8)
4.3.4 Aggregation strategy
By default, we adopt the mean aggregation strategy to obtain explicit point representations. We also investigate two alternative strategies: max strategy, which selects the maximum pixel representation across different views as the corresponding explicit point representation; network-parametric strategy, which employs an MLP followed by a softmax layer to compute aggregation weights of the pixel representations. As shown in Table V, PointAD+ is not restricted to a specific aggregation scheme but generalizes across different strategies. We adopt the mean strategy as it provides a favorable trade-off between performance and computational overhead.
@@ 4.3.5 Complexity analysis
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table VI, including GFLOPs, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. PointAD+ requires only about 2.0 more GFLOPs than PointAD (1658.34 vs. 1656.34) yet achieves a significant performance improvement. In contrast, other CLIP-based methods, such as PointCLIP V2, consume larger GFLOPs but deliver worse results. Although CLIP+Rendering, Cheraghian, and AnomalyCLIP use fewer GFLOPs, they fail to provide competitive performance. As for inference speed, compared to PointCLIP V2, PointAD requires less time to infer an image and achieves a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and multimodal 3D. CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. From the above results, PointAD+ achieves a favorable trade-off between performance and overhead.
TABLE VII: Ablation on proposed modules. F stage and S stage represent the first and second stages of G-aggregation, respectively.
$L^{r}$ $L^{g}$ $L^{c}$ Point detection Multimodal detection
@@
TABLE VIII: Ablation on the neighbors $k$.

### v5

@@ 3.6.1 Implicit point representation learning on the rendering layer (PointAD)
$L_{I(3d)}^{global}=\frac{1}{N}\sum_{i}\textrm{CE}(\frac{1}{K}\textstyle\sum_{f_{i}^{(k)}\in p_{i}}P(g^{r}_{c},f_{i}^{(k)}),\max{(y^{3d}_{i})}).$ (6)
@@ 3.6.3 Cross-hierarchy contrastive alignment
$-\log\frac{\exp(\cos(g^{g}_{n},g^{r}_{n}))}{\exp(\cos(g^{g}_{n},g^{r}_{a}))+\exp(\cos(g^{g}_{n},g^{r}_{n}))}$ (18)
$-\log\frac{\exp(\cos(g^{g}_{a},g^{r}_{a}))}{\exp(\cos(g^{g}_{a},g^{r}_{a}))+\exp(\cos(g^{g}_{n},g^{r}_{a}))}$
$-\log\frac{\exp(\cos(g^{g}_{n},g^{r}_{n}))}{\exp(\cos(g^{g}_{a},g^{r}_{n}))+\exp(\cos(g^{g}_{n},g^{r}_{n}))}$
@@ 4.1 Experiment Setup
Dataset. We evaluate the performance of ZS 3D anomaly detection on three publicly available 3D anomaly detection datasets, MVTec3D-AD, Eyecandies, and Real3D-AD. MVTec3D-AD comprises 4147 point clouds across 10 categories. These objects exhibit diverse object semantics, including bagel, cable gland, carrot, cookie, dowel, foam, peach, potato, rope, and tire. The training dataset comprises 2656 normal point clouds, and the test dataset includes 948 normal and 249 anomaly point clouds, covering several anomaly types. Point-wise annotations are available for the point clouds. MVTec3D-AD also provides corresponding 2D-RGB image counterparts for the point clouds. We remove the background plane of point clouds in the whole dataset like [horwitz2023back]. Eyecandies also has 10 different classes and provides the corresponding 2D RGB information. The more recent Real3D-AD is a challenging dataset containing stereo point clouds with 12 categories, without the corresponding RGB information.
@@
Implementation details. Both point clouds and 2D renderings are resized to 336 $\times$ 336. We use Open3D library to generate 9 views by rotating point clouds along the X-axis at angles of $\{-\frac{4}{5}\pi,-\frac{3}{5}\pi,-\frac{2}{5}\pi,-\frac{1}{5}\pi,0,\frac{1}{5}\pi,\frac{2}{5}\pi,\frac{3}{5}\pi,\frac{4}{5}\pi\}$ for most categories. Some objects lose their surface when rendered at certain angles, we manually adjust them to an appropriate angle. The backbone of PointAD+ and PointAD is the pre-trained CLIP model (VIT-L/14@336px in open_clip). Following [zhou2023anomalyclip], we improve the local visual semantics of the vision encoder of CLIP without modifying its parameters. During training, we keep all parameters of CLIP frozen and set the learnable word embeddings in object-agnostic text templates to $12$. The hyperparameters, including the number of neighbors $k$, the weighting scalar $\alpha$, and the scale parameter $\sigma$ of the Gaussian kernel, are set to 10, 0.5, and 1, respectively. We use the Adam optimizer with a learning rate of 0.001 to optimize the learnable parameters. The experiment runs for 15 epochs with a batch size of 4. All experiments were conducted on a single NVIDIA RTX 3090 24GB GPU using PyTorch-2.0.0. PointAD+ and PointAD by default use object-agnostic text prompts, whereas PointAD+-CoOp and PointAD-CoOp employ object-aware prompts.
Baselines. As ZS 3D and multimodal 3D anomaly detection remain underexplored areas in the field, we make a great effort to provide these comparisons. We apply the original CLIP to our framework for 3D detection, called CLIP + Rendering. Also, we reproduce SOTA 3D recognition works including PointCLIP V2 [zhu2023pointclip] and Cheraghian [cheraghian2022zero] and adapt them for ZS 3D anomaly detection. We compare the SOTA 2D anomaly detection approach AnomalyCLIP [zhou2023anomalyclip] by fine-tuning it on depth maps. All hyperparameters in these baselines are kept the same. We will present the detailed reproduction:
  • CLIP + Rendering is a method, where we apply the original CLIP into our framework for ZS 3D anomaly detection. It uses the same rendering procedure as PointAD+ and PointAD. Following [jeong2023winclip, zhou2023anomalyclip], we integrate anomaly semantics into CLIP by two class text prompt templates: A photo of a normal [cls] and A photo of an anomalous [cls], where cls denotes the target class name.
@@
  • Cheraghian (IJCV 2022) is an approach for ZS 3D classification without foundation models. They directly extract the point presentations by PointNet and use word2vector [mikolov2013efficient] to generate the textual embedding of [c]. To incorporate the anomaly semantics into Cheraghian, we average the textual embeddings of [c] and damaged. We replace the global representation with dense representations to provide the segmentation.
@@ 4.2 Main Results
ZS 3D anomaly detection. Table I presents the comparison of ZS 3D performance. Compared to the point-based method Cheraghian and the projection-based method PointCLIP V2, PointAD+ and PointAD achieve superior performance on ZS 3D anomaly detection over all three datasets across 32 object semantics. Especially, PointAD outperforms CLIP + Rendering from 61.2% to 82.0% I-AUROC and from 85.8% to 94.2% AP on MVTec3D-AD. PointAD+ further obtains the performance gain over PointAD. Besides global anomaly detection, PointAD+ and PointAD also attain superior segmentation performance. PointAD improves MVTec3D-AD by a large margin compared to Cheraghian from 88.2% to 95.5% P-AUROC and from 57.0% to 84.4% AUPRO. This improvement in overall performance is attributed to PointAD adapting CLIP’s strong generalization to glocal anomaly semantics through hybrid representation learning. On this basis, PointAD+ further improves the performance from 95.5% to 96.8% P-AUROC and from 84.4% to 88.4% AUPRO. It is attributed to the hierarchy representation learning, incorporating explicit point relations to better reveal the abnormality within point clouds. As shown in Figure 5, PointAD+ and PointAD could accurately segment the unseen anomalies across diverse objects. In addition, PointAD+ and PointAD advance PointAD+-CoOp and PointAD-CoOp across all datasets by blocking the class semantics in text prompts [zhou2023anomalyclip].
ZS multimodal 3D anomaly detection. PointAD+ and PointAD allow the direct incorporation of RGB counterpart when it is available during the test. Table II indicates that PointAD+ and PointAD leverage additional RGB information in a plug-and-plug manner to boost their performance. PointAD gets the performance gain from 82.0% to 86.9% AUROC and from 94.2% to 96.1% AP for global semantics on MVTec3D-AD. As for local semantics, the performance improves from 95.5% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. A large performance gain is also obtained on Eyecandies and Real3D-AD. PointAD+ can also benefit from the RGB information and break the performance achieved by PointAD, i.e., 98.2% vs. 97.2% P-AUROC and 93.1% vs. 90.2% AUPRO on MVTec3D-AD, and 79.2% vs. 77.7% I-AUROC and 81.7% vs. 80.4% AP on Eyecandies. As other methods can improve their performance in some metrics, they still suffer from performance degradation in other metrics due to inefficient integration of the two modalities. Instead, PointAD+ achieves overall improvement by incorporating explicit joint constraints on both point and pixel information.
@@
4.3 Result Analysis
PointAD+ characterizes itself by capturing abnormality from two aspects: implicit 3D abnormality, where point anomalies are projected into the pixel space; explicit 3D abnormality, which reveals the inherent point relations. In this section, we provide a more in-depth analysis of how they contribute to the success of PointAD+.
@@ 4.3 Result Analysis
Figure 8: Visualization on stereo point clouds (Real3D-AD)
@@
4.3.2 Implicit point learning on rendering layer
We first investigate how rendering anomalies facilitate 3D anomaly detection. As illustrated in Figure 7, we visualize the anomaly score maps of the 3D and corresponding 2D counterparts of PointAD+ on MVTec3D-AD. And, we also visualize the anomaly score map of 2D counterparts according to Eq. 9, projected by 3D point anomalies. It can be observed that PointAD+ has a strong detection ability on such 2D anomalies. The strong representative pixel representations from multiple views facilitate more precise 3D anomaly detection. We conduct a quantitative analysis of the view number in the ablation study. The strong 3D and 2D detection capabilities of PointAD+ are from hybrid representation learning, which not only enables PointAD+ to capture the 3D anomalies but also explicitly constrains 2D representations.
4.3.3 Explicit point learning on geometric layer
We further analyze what PointAD+ learns from the geometry layer as defined in Eq. 15. As illustrated in Figure 8, the explicit point learning module excels at capturing abnormalities in stereo point clouds, as their abnormalities rely heavily on underlying spatial relations. By incorporating geometric information, it exhibits a stronger capability to detect abnormal point relations, outperforming implicit point learning in identifying such structural anomalies. However, explicit point learning may introduce false positives near object boundaries in the point cloud. This observation is consistent with our discussion in Section 3.3.
TABLE V: Comparison of computation overhead with SOTA approaches on MVTec3D-AD.
@@ 4.3.3 Explicit point learning on geometric layer
PointCLIP V2 4660.49 0.66 9747MB (87.4, 52.3) (78.3, 49.4)
CLIP + Rendering 1057.36 3.61 3685MB (61.2, 85.8) (-, 54.4)
Cheraghian 31.49 2.86 4847MB (88.2, 57.0) (53.6, 81.7)
AnomalyCLIP 1656.34 5.26 3348MB (88.9, 60.9) (56.4, 83.5)
PointAD 1656.34 2.52 4275MB (95.5, 84.4) (82.0, 94.2)
PointAD+ 1658.34 2.18 5811MB (96.9, 88.7) (82.5, 94.8)
@@ 4.3.4 Complexity analysis
The computation overhead is also an indicator to assess the effectiveness of the proposed method. We compare the computation overhead in Table V, including GFLOPs, frames per second (FPS), and GPU memory consumption with a batch size of 1. For a fair comparison, we maintain NVIDIA RTX 3090 24GB GPU idle. PointAD+ requires only about 2.0 more GFLOPs than PointAD (1658.34 vs. 1656.34) yet achieves a significant performance improvement. In contrast, other CLIP-based methods, such as PointCLIP V2, consume larger GFLOPs but deliver worse results. Although CLIP+Rendering, Cheraghian, and AnomalyCLIP use fewer GFLOPs, they fail to provide competitive performance. As for inference speed, compared to PointCLIP V2, PointAD requires less time to infer an image and achieves a higher 2.52 vs. 0.66 FPS with lower graphic memory usage. PointAD+ needs a slightly higher computation overhead over PointAD, i.e., 2.52 vs. 2.18 FPS, and attains a noticeable performance improvement on ZS 3D and multimodal 3D. CLIP + rendering has the fastest inference speed, but its detection performance lags behind PointAD+ and PointAD. From the above results, PointAD+ achieves a favorable trade-off between performance and overhead.
5 Ablation Study
TABLE VI: Ablation on proposed modules. F stage and S stage represent the first and second stages of G-aggregation, respectively.
$L_{hybrid}^{r}$ $L^{geo}$ $L^{c}$ Point detection Multimodal detection
@@ 5 Ablation Study
5.1 Ablation on modules and key hyperparameters
Module ablation. We evaluate the effectiveness of the proposed techniques by progressively incorporating the introduced modules. As shown in Table VI, the baseline vanilla model, which represents CLIP + rendering, performs poorly on both 3D and multimodal 3D anomaly detection due to CLIP’s focus on aligning 2D object semantics rather than anomaly semantics. Adding the 3D global branch incorporates global anomaly semantics into PointAD+, improving both local and global anomaly detection performance. Although the 3D local branch further enhances performance, local accuracy on multimodal 3D detection declines due to the absence of 2D constraints, leading to inefficient multimodal fusion when integrating 2D RGB information in a plug-and-play manner. The inclusion of the 2D global branch explicitly introduces 2D anomaly information, enabling PointAD+ to achieve overall performance gains and better multimodal fusion. By focusing on 2D anomaly regions, PointAD+ develops a deeper understanding of point clouds from 2D representations, improving multimodal segmentation performance from 92.9% to 97.2% P-AUROC and from 84.4% to 90.2% AUPRO. So far, PointAD+ does not explicitly account for point relationships within point clouds. To address this, explicit point learning is introduced to model abnormal spatial relationships. PointAD+ leverages geometric characteristics, achieving notable improvements in segmentation performance. Specifically, P-AUROC increases from 95.5% to 96.3% and AUPRO rises from 84.4% to 87.5% on ZS 3D detection. On multimodal 3D detection, P-AUROC improves from 97.2% to 97.9% and AUPRO increases from 90.2% to 92.5%. Finally, cross-alignment is introduced to enhance interactions between hierarchical representations, leading to further performance gains over these metrics.
Ablation on the number of neighbors The number of neighbors $k$ determines the receptive field of each point within its local neighborhood. As illustrated in Table VII, we further explore $k\in\{5,10,15,20,25\}$ to study its impact. For point detection, the best local accuracy is achieved at $k=10$ (96.9 AUROC and 88.7 AUPRO), and the best global accuracy is also observed at $k=10$ (82.5 AUROC and 94.8 AP). For multimodal detection, both $k=10$ and $k=20$ yield superior performance. Overall, the detection results remain stable across different values of $k$. This indicates that PointAD+ is robust to the number of neighboring points.
TABLE VII: Ablation on the neighbors $k$.